OpenStudy (anonymous):

The midpoint of line segment from p1to p2 is (-7,4). If p1=(-5,6). What is p2?

OpenStudy (mertsj):

\[\frac{-5+x}{2}=-7\]

OpenStudy (mertsj):

\[\frac{6+y}{2}=4\]

OpenStudy (anonymous):

So which is the answer, if you don't mind me asking?

OpenStudy (anonymous):

@ Mertsj, thanks for the response, which is the answer?

OpenStudy (mertsj):

What does the problem ask you to find?

OpenStudy (whpalmer4):

@Redbone12 you give us your answer, and we'll tell you if you got it correct...